# Can We Trust AI Benchmarks? An Interdisciplinary Review of Current Issues in AI Evaluation

### v1

@@ 4. METHODOLOGY
In the process, we surveyed a wide range of papers that, for instance, highlight general problems with the production of AI datasets (Mitchell et al., 2019; Orr and Crawford, 2024a), propose alternative ways of designing benchmark leaderboards (Rodriguez et al., 2021; Liu et al., 2021), or discuss the wider use of proxies in tests and evaluations (Mulvin, 2021; Pinch, 1993; Marres and Stark, 2020). Such papers provide important contextual insights to discussions concerning benchmarks. However, they were omitted from what we considered our core collection of previous research. This collection ended up consisting of about 100 papers and articles that explicitly and primarily highlight issues with benchmarks. Notably, articles that propose new benchmarks were not added to this collection by default, even though such articles naturally contain some level of benchmark critique. The reasoning behind this is twofold. On the one hand, it was necessary to limit the size and scope of our review. On the other hand, many articles proposing new benchmarks (if not most of them) do not question or discuss many (if any) of the underlying assumptions that benchmark assessments rely on. In that sense, they reproduce the general notion that quantitative benchmarks provide a reasonable technical ”fix” to issues with AI safety and capability assessments, rather than taking part in more fundamental and critical discussions on current limitations in benchmark use and design. For these reasons, they were also excluded from our core collection of benchmark critique.
Our collection was further limited by only containing articles published between 1st January 2014 and 31st December 2024. While it is true that AI models have been tested and evaluated since their origins in the mid 1950’s (with the Turing test serving as an iconic example), we limit our review to this decade long period, since 2014 marks the starting point for recent intensifications in AI research and development. We consider both published and pre-published papers in our review, which on a practical level involved categorising identified critique into relevant areas of concern, without a pre-existing target of final categories. When important concepts (such as sandbagging) appeared in the literature, targeted searches for research in these fields were made. Given that a majority of existing AI benchmarks (and benchmark critique) address text-based AI models, we also made special efforts to include discussions concerning benchmarks for images, sound, and audiovisual content.
Our resulting meta-review is not exhaustive but it covers a broad range of critique that has been aimed at benchmarking practices. In the following discussions, we focus on works that voice critique with particular relevance for policy makers and policy implementation. We also shed light on works that emphasise areas of concern that cut across different modalities (text, images, sound, moving images) and point toward fundamental weaknesses in the design and application of benchmarks, as opposed to research that critiques individual benchmarks. Our goal has further been to provide a diverse account of concerns regarding benchmarks. Hence, some sections will have ample publications reinforcing their claims while others, that voice more recent and/or unique critique, are backed up by fewer sources.
5. NINE REASONS TO BE CAUTIOUS WITH BENCHMARKS
@@
5.1. Problems with Data Collection, Annotation, and Documentation
@@
An initial set of issues with AI benchmarks found during our research is limitations in the collection, annotation, and documentation of benchmark datasets. This ties into a broader critique regarding insufficient documentation in AI research which is central to calls for more transparent and trustworthy algorithmic systems (Gebru et al., 2021; Mitchell et al., 2019; Orr and Crawford, 2024a; Simson et al., 2024; Scheuerman et al., 2021). Research has found that it is often difficult to trace precisely how, when, and by whom benchmark datasets have been made (Reuel et al., 2024; Denton et al., 2020) which compromises the ability for benchmarks to be used in robust and generalisable ways (Arzt and Hanbury, 2024). The issue has partly been linked to the low status of dataset-related work within the machine learning community (which instead privileges model development) (Orr and Crawford, 2024b; Sambasivan et al., 2021), and the fact that AI datasets are often “reduced, reused, and recycled” (Koch et al., 2021), which complicates documentations of their possible limitations (Thylstrup et al., 2022; Park and Jeoung, 2022). Notably, Koch et al. (2021) have found that more than 70 percent of the benchmark datasets used in prominent computer vision papers had been reused from other domains. Park and Jeoung (2022) also find that while benchmark datasets often contain ample information on how to use the dataset, documentation is often missing concerning their shortcomings and social impact. In their exploration of benchmark sharing platforms like HuggingFace [^1] https://huggingface.co/. Last seen 22nd January 2025. and PapersWithCode [^2] https://paperswithcode.com/. Last seen 22nd January 2025. the same researchers also note that confusing metadata terminology severely complicates efforts to understand benchmark dataset documentation.
@@
Like AI training datasets, benchmarks have been singled out for raising ethical and legal questions concerning copyrights, privacy, informed consent, and rights to opt-out (Paullada et al., 2021). Many benchmarks rely on crowd-sourced or user-generated content from platforms like Wikihow [^3] https://www.wikihow.com/Main-Page. Last seen 22nd January 2025., Reddit [^4] https://www.reddit.com/. Last seen 22nd January 2025., or trivia websites (Keegan, 2024; Grill, 2024), whose annotation may be noisy and biased, lack input from expertise in specialised fields, or be produced under exploitative conditions (Tsipras et al., 2020; Aroyo and Welty, 2015; Sen et al., 2015). Previous research has noted that this matters when baselines for ”good”, ”bad”, and ”safe enough” AI models are calibrated in sensitive contexts (Rauh et al., 2024; Grill, 2024). Moreover, it is essential to highlight the absence of human performance references and difficulty rubrics in benchmarks, which are increasingly important factors in evaluating capabilities and generality (Chollet, 2019).
@@
A lack of care in the making of benchmark datasets has further been found to result in AI models exploiting unknown quirks and spurious cues in the training data, rather than solving their original intended task (Liao et al., 2021; Paullada et al., 2021; Geirhos et al., 2020).For instance, Oakden-Rayner et al. (2019) found that an X-ray image classification model predicted collapsed lungs with high accuracy, although it turned out the model only identified the presence of a chest drain (used to cure the condition) that was (unknowingly) present in a majority of the positive training images. In other words, the model performed well in benchmark tests, but the reason why completely sidestepped the original purpose of the task (identifying collapsed lungs). When the all images containing chest drains were removed from the training dataset, the model performance dropped by over 20%. Similar patterns have also been confirmed in recent research on LLM evaluation (Pacchiardi et al., 2024) and is sometimes discussed as failure in the ”world model” of an AI system (Vafa et al., 2024). While it could be argued that picking up on spurious cues is a sign of intelligence and therefore a capability worth noting, the issues above highlight how a lack of in-depth attention to the data that benchmarks rely on - alongside a failure to acknowledge how little is actually known about how and why AI models perform well in benchmark tests - can produce frail and uncertain AI evaluations.
5.2. Weak Construct Validity and Epistemological Claims
@@
Another genre of benchmark critique focuses on the epistemological claims that tend to surround benchmarks and examines the limits of what can be known through quantitative AI tests. A central reference point in these discussions is the observation by Raji et al. (2021) that many benchmarks suffer from construct validity issues in the sense that they do not measure what they claim to measure. As the authors proclaim, this is especially troublesome when benchmarks promise to measure universal or general capabilities, since this vastly misrepresents their actual capability. As a result, the authors argue that framing an AI benchmark dataset as general purpose ”is ultimately dangerous and deceptive, resulting in misguidance on task design and focus, underreporting of the many biases and subjective interpretations inherent in the data as well as enabling, through false presentations of performance, potential model misuse”. (Raji et al., 2021, p. 5). At the heart of this critique lies the realization that many benchmarks do not have a clear definition of what they claim to measure, which makes it impossible to measure if they succeed in the task or not (Blodgett et al., 2021; Bartz-Beielstein et al., 2020). In a close analysis of four benchmarks used to evaluate fairness in natural language processing (StereoSet, CrowS-Pairs, WinoBias, and WinoGender), Blodgett et al. (2021) for example found that all benchmarks revealed severe weaknesses in terms of defining what is being measured. For instance, culturally complex and highly contested concepts like ”stereotypes” or ”offensive language” were left unspecified, causing a series of logical failures and interpretational conflicts. Elsewhere, research has shown strong disagreements in how benchmark tasks are conceptualised and operationalised (Subramonian et al., 2023), and found that benchmarks are applied in highly idiosyncratic ways (Röttger et al., 2024). Frequently, the difficulty defining what benchmarks evaluate persist since there is no clear, stable and absolute ground truth for what is claimed to be measured (Narayanan and Kapoor, 2023b). Since concepts like ”bias” and ”fairness” are inherently contested, messy, and shifting, benchmarks that promise to measure such terms will inevitably suffer from an ”abstraction error” that produces a false sense of certainty (Selbst et al., 2019, p. 63).
@@
It has also been pointed out that many benchmarks datasets are inadequate and/or unuseful proxys for what they are meant to evaluate. For instance, researchers have identified a slippage in distinguishing between algorithmic ”harms” and algorithmic ”wrongs” when evaluating the capabilities of AI models - two not necessarily overlapping concepts (Diberardino et al., 2024). Others have questioned whether the content of benchmark datasets are reasonable substitutes for the ”real world” scenarios they are meant to reflect. For instance, researchers have questioned the decision to use examples generated by Amazon crowdworkers and posts from the Reddit forum ”Am I the asshole?” as proxies for ethics and morals in benchmarks such as HellaSwag - a widely cited benchmark for language models (Keegan, 2024). In benchmarks consisting of professional exams, researchers have further argued that such tests ”emphasize the wrong thing” and ”overemphasize precisely the thing that language models are good at” and are thus unreliable measures of things such as medical or legal skill (Narayanan and Kapoor, 2023b, n.p). As Narayanan and Kapoor (2023b) put it, ”it’s not like a lawyer’s job is to answer bar exam questions all day”. A recent study by Ren et al. (2024) also found that many widely used safety benchmarks (including ETHICS, TruthfulQA, GPQA, QuALITY, MT-Bench, LMSYS Chatbot ARENA, ANLI, AdvGLUE, and AdvGLUE++) highly correlate with general and upstream model capabilities, raising concerns regarding “safetywashing” as applying them could imply that “capability improvements are misrepresented as safety advancements”. While it could be argued that capability and safety are largely entwined (the more capable a model is, the higher the likelihood it could cause harm), Ren et al. (2024, p. 1) suggest that a blurred distinction between the two may hide the fact that severe biases and safety issues in AI models can persist, even as their overall capabilities improve.
5.3. Sociocultural Context and Gap
@@
Another key insight from previous research concerns the importance of the social, economic and cultural contexts where AI benchmarks are created, used, and maintained. Among researchers engaging in benchmark critique, we identify a strong consensus regarding the need to recognise that benchmarks are ultimately ”normative instruments that perpetuate particular epistemological perspectives about how the world is ordered” (Orr and Kang, 2024, p. 1877). Qualitative research has also examined the cultural and social environments where benchmarks are made, finding that they are deeply shaped by shared and arbitrary assumptions, commitments, and dependencies (Engdahl, 2024; Michael et al., 2022; Scheuerman et al., 2021; Orr and Crawford, 2024b; Sambasivan et al., 2021; Paullada et al., 2021). Such assumptions for example include valuing ”efficiency at the expense of care; universality at the expense of contextuality; impartiality at the expense of positionality; and model work at the expense of data work” (Scheuerman et al., 2021), or reproduce contested ideas such as the notion that low-quality data can be drowned out by scale (Orr and Crawford, 2024b). Recent research has also illustrated that AI safety research and benchmark competitions are increasingly informed by political movements and ideologies such as longtermism and effective altruism (Ahmed et al., 2024)
@@
Scholars have further identified sociotechnical gaps and a lack of consideration for downstream utility as a concern in AI benchmarking (Hutchinson et al., 2022). Liao and Xiao (2023) highlight that this means it is often unclear who is meant to care about benchmark evaluation results and how they should be used in practice. For instance, a recent study by Blagec et al. (2023), which compared the explicitly stated needs for AI technologies among clinical medical practitioners with existing clinical benchmark datasets, found that most benchmarks failed to answer to the needs of medical experts. They also found that benchmarks for the most urgently requested medical or clinical tasks were completely missing and noted that similar misalignments likely exist in many other fields. For example, Jannach and Bauer (2020, p. 79) highlight a lack of attention to how recommender systems ”may create value; how they, positively or negatively, impact consumers, businesses, and the society; and how we can measure the resulting effects”. Ethayarajh and Jurafsky (2021) further argue that failures to consider the practical utility of benchmarks has for example made it possible to ignore the discriminatory and environmental damages of AI technologies and allowed for highly energy-inefficient and deeply biased AI models to reach the top of most benchmark leaderboards.
5.4. Narrow Benchmark Diversity and Scope

### v2

@@ 4 Methodology
In the process, we surveyed a wide range of papers that, for instance, highlight general problems with the production of AI datasets [Mitchell et al., 2019; Orr and Crawford, 2024b], propose alternative ways of designing benchmark leaderboards [Rodriguez et al., 2021; Liu et al., 2021], or discuss the wider use of proxies in tests and evaluations [Mulvin, 2021; Pinch, 1993; Marres and Stark, 2020]. Such papers provide important contextual insights to discussions concerning benchmarks. However, they were omitted from what we considered our core collection of previous research. This collection ended up consisting of about 110 papers and articles that explicitly and primarily highlight issues with benchmarks. Notably, articles that propose new benchmarks were not added to this collection by default, even though such articles naturally contain some level of benchmark critique. The reasoning behind this is twofold. On the one hand, it was necessary to limit the size and scope of our review. On the other hand, many articles proposing new benchmarks (if not most of them) do not question or discuss many (if any) of the underlying assumptions that benchmark assessments rely on. In that sense, they reproduce the general notion that quantitative benchmarks provide a reasonable technical "fix" to issues with AI safety and capability assessments, rather than raising more fundamental and critical discussions on limitations in benchmark use and design. For these reasons, they were also excluded from our core collection of papers. Our collection was further limited by only containing articles published between 1st January 2014 and 31st December 2024. While it is true that AI models have been tested and evaluated since their origins in the mid 1950’s (with the Turing test serving as an iconic example), we limit our review to this decade long period, since 2014 marks the starting point for recent intensifications in AI research and development.
Our resulting meta-review is not exhaustive but it covers a broad range of critique that has been aimed at benchmarking practices. The nine issue categories presented were identified after close-reading and relevant articles, classifying and grouping papers voicing similar critique, and discussing these classifications within the wider author group. The resulting issue areas represent the result of these discussions, although many issues were difficult to categorize. E.g., problems with Anglo-centric benchmarking datasets could simultaneously be framed as a cultural problem (Anglo-centrism), a construct validity problem (as English is taken to represent culture as a whole), and a problem with narrow benchmark diversity and scope. As a result of such overlaps, identified issues are not meant to be read as an absolute and all-encompassing definition of benchmark vulnerabilities, but as a narrative tool to present our findings. We especially focus on works that voice critique with relevance for policy makers and policy implementation, highlight areas of concern that cut across different modalities (text, images, sound, moving images), and point toward fundamental weaknesses in the design and application of benchmarks, as opposed to research that critiques individual benchmarks. Our goal has further been to provide a diverse account of concerns regarding benchmarks.
5 Nine Reasons to Be Cautious with Benchmarks
@@
5.1 Problems with Data Collection, Annotation, and Documentation
@@
An initial set of issues with AI benchmarks found during our research is limitations in the collection, annotation, and documentation of benchmark datasets. This ties into a broader critique regarding insufficient documentation in AI research which is central to calls for more transparent and trustworthy algorithmic systems [Gebru et al., 2021; Mitchell et al., 2019; Orr and Crawford, 2024b; Simson et al., 2024; Scheuerman et al., 2021]. Research has found that it is often difficult to trace precisely how, when, and by whom benchmark datasets have been made [Reuel et al., 2024; Denton et al., 2020] which compromises the ability for benchmarks to be used in robust and generalisable ways [Arzt and Hanbury, 2024]. The issue has partly been linked to the low status of dataset-related work within the machine learning community (which instead privileges model development) [Orr and Crawford, 2024a; Sambasivan et al., 2021], and the fact that AI datasets are often “reduced, reused, and recycled” [Koch et al., 2021], which complicates documentations of their possible limitations [Thylstrup et al., 2022; Park and Jeoung, 2022]. Notably, Koch et al. [2021] have found that more than 70 percent of the benchmark datasets used in prominent computer vision papers had been reused from other domains. Park and Jeoung [2022] also find that while benchmark datasets often contain ample information on how to use the dataset, documentation is often missing concerning their shortcomings and social impact. In their exploration of benchmark sharing platforms like HuggingFace [^1] https://huggingface.co/. Last seen May 25, 2025. and PapersWithCode [^2] https://paperswithcode.com/. Last seen May 25, 2025. the same researchers also note that confusing metadata terminology severely complicates efforts to understand benchmark dataset documentation.
@@
Like AI training datasets, benchmarks have been singled out for raising ethical and legal questions concerning copyrights, privacy, informed consent, and rights to opt-out [Paullada et al., 2021]. Many benchmarks rely on crowd-sourced or user-generated content from platforms like Wikihow [^3] https://www.wikihow.com/. Last seen May 25, 2025., Reddit [^4] https://www.reddit.com/. Last seen May 25, 2025., or trivia websites [Keegan, 2024; Grill, 2024], whose annotation may be biased, lack input from expertise in specialised fields, or be produced under exploitative conditions [Tsipras et al., 2020; Aroyo and Welty, 2015; Sen et al., 2015]. Previous research has noted that this matters when baselines for "good", "bad", and "safe enough" AI models are calibrated in sensitive contexts [Rauh et al., 2024; Grill, 2024]. Moreover, the absence of human performance references and difficulty rubrics in benchmarks has been highlighted, which are increasingly considered important factors in evaluating capabilities and generality [Chollet, 2019].
@@
Noisy human annotations and lack of care in the making of benchmark datasets have further been found to skew benchmark scores [Kejriwal et al., 2024] and result in AI models exploiting unknown quirks and spurious cues in the training data, rather than solving their original intended task [Liao et al., 2021; Paullada et al., 2021; Geirhos et al., 2020]. For instance, Oakden-Rayner et al. [2019] found that an X-ray image classification model predicted collapsed lungs with high accuracy, although it turned out the model only identified the presence of a chest drain (used to cure the condition) that was (unknowingly) present in a majority of the positive training images. In other words, the model performed well in benchmark tests, but the reason why completely sidestepped the original purpose of the task (identifying collapsed lungs). When the all images containing chest drains were removed from the training dataset, the model performance dropped by over 20%. Similar patterns have also been confirmed in recent research on LLM evaluation [Pacchiardi et al., 2024] and is sometimes discussed as failure in the "world model" of an AI system [Vafa et al., 2024]. While it could be argued that picking up on spurious cues is a sign of intelligence and therefore a capability worth noting, the issues above highlight how a lack of in-depth attention to the data that benchmarks rely on - alongside a failure to acknowledge how little is actually known about how and why AI models perform well in benchmark tests - can produce frail and uncertain AI evaluations.
5.2 Weak Construct Validity and Epistemological Claims
@@
Another genre of benchmark critique focuses on the epistemological claims that tend to surround benchmarks and examines the limits of what can be known through quantitative AI tests. A central reference point in these discussions is the observation by Raji et al. [2021] that many benchmarks suffer from construct validity issues in the sense that they do not measure what they claim to measure. As the authors proclaim, this is especially troublesome when benchmarks promise to measure universal or general capabilities, since this vastly misrepresents their actual capability. As a result, the authors argue that framing an AI benchmark dataset as general purpose "is ultimately dangerous and deceptive, resulting in misguidance on task design and focus, underreporting of the many biases and subjective interpretations inherent in the data as well as enabling, through false presentations of performance, potential model misuse". [Raji et al., 2021, p. 5]. At the heart of this critique lies the realization that many benchmarks do not have a clear definition of what they claim to measure, which makes it impossible to measure if they succeed in the task or not [Blodgett et al., 2021; Bartz-Beielstein et al., 2020]. In a close analysis of four benchmarks used to evaluate fairness in natural language processing (StereoSet, CrowS-Pairs, WinoBias, and WinoGender), Blodgett et al. [2021] for example found that all benchmarks revealed severe weaknesses in terms of defining what is being measured. For instance, culturally complex and highly contested concepts like "stereotypes" or "offensive language" were left unspecified, causing a series of logical failures and interpretational conflicts. Elsewhere, research has shown strong disagreements in how benchmark tasks are conceptualised and operationalised [Subramonian et al., 2023], and found that benchmarks are applied in highly idiosyncratic ways [Röttger et al., 2024] that can be misleading in numerous ways [Leech et al., 2024]. Frequently, the difficulty in defining what benchmarks evaluate persists since there is no clear, stable, and absolute ground truth for what is claimed to be measured [Narayanan and Kapoor, 2023a]. Since concepts like "bias" and "fairness" are inherently contested, messy, and shifting, benchmarks that promise to measure such terms will inevitably suffer from an "abstraction error" that produces a false sense of certainty [Selbst et al., 2019, p. 63].
@@
It has also been pointed out that many benchmark datasets are inadequate and/or useless proxies for what they are meant to evaluate. For instance, researchers have identified a slippage in distinguishing between algorithmic "harms" and algorithmic "wrongs" when evaluating the capabilities of AI models - two not necessarily overlapping concepts [Diberardino et al., 2024]. Others have questioned whether the content of benchmark datasets is a reasonable substitute for the "real world" scenarios they are meant to reflect. For instance, the decision to use examples generated by Amazon crowdworkers and posts from the Reddit forum "Am I the asshole?" as proxies for ethics and morals in benchmarks such as HellaSwag - a widely cited benchmark for language models has been questioned [Keegan, 2024]. In benchmarks consisting of professional exams, researchers have further argued that such tests "emphasize the wrong thing" and "overemphasize precisely the thing that language models are good at" and are thus unreliable measures of things such as medical or legal skill [Narayanan and Kapoor, 2023a, n.p]. As Narayanan and Kapoor [2023a] put it, "it’s not like a lawyer’s job is to answer bar exam questions all day". A recent study by Ren et al. [2024] also found that many widely used safety benchmarks (including ETHICS, TruthfulQA, GPQA, QuALITY, MT-Bench, LMSYS Chatbot ARENA, ANLI, AdvGLUE, and AdvGLUE++) highly correlate with general and upstream model capabilities, raising concerns regarding “safetywashing” as applying them could imply that “capability improvements are misrepresented as safety advancements". While it could be argued that capability and safety are largely entwined (the more capable a model is, the higher the likelihood it could cause harm), Ren et al. [2024, p. 1] suggest that a blurred distinction between the two may hide the fact that severe biases and safety issues in AI models can persist, even as their overall capabilities improve.
5.3 Sociocultural Context and Gap
@@
Another key insight from previous research concerns the importance of the social, economic and cultural contexts where AI benchmarks are created, used, and maintained. Among researchers engaging in benchmark critique, we identify a strong consensus regarding the need to recognise that benchmarks are ultimately "normative instruments that perpetuate particular epistemological perspectives about how the world is ordered" [Orr and Kang, 2024, p. 1877]. Qualitative research has also examined the cultural and social environments where benchmarks are made, finding that they are deeply shaped by shared and arbitrary assumptions, commitments, and dependencies [Engdahl, 2024; Michael et al., 2022; Scheuerman et al., 2021; Orr and Crawford, 2024a; Sambasivan et al., 2021; Paullada et al., 2021]. Such assumptions for example include valuing "efficiency at the expense of care; universality at the expense of contextuality; impartiality at the expense of positionality; and model work at the expense of data work" [Scheuerman et al., 2021], or reproduce contested ideas such as the notion that low-quality data can be drowned out by scale [Orr and Crawford, 2024a]. Recent research also illustrated that AI safety research and benchmark competitions are increasingly informed by political movements and ideologies such as longtermism and effective altruism [Ahmed et al., 2024].
@@
Scholars have further identified sociotechnical gaps and a lack of consideration for downstream utility as a concern in AI benchmarking [Hutchinson et al., 2022; Frieder et al., 2024]. Liao and Xiao [2023] highlight that this means it is often unclear who is meant to care about benchmark evaluation results and how they should be used in practice. For instance, a recent study by Blagec et al. [2023], which compared the explicitly stated needs for AI technologies among clinical medical practitioners with existing clinical benchmark datasets, found that most benchmarks failed to answer to the needs of medical experts. They also found that benchmarks for the most urgently requested medical or clinical tasks were completely missing and noted that similar misalignments likely exist in many other fields. For example, Jannach and Bauer [2020, p. 79] highlight a lack of attention to how recommender systems "may create value; how they, positively or negatively, impact consumers, businesses, and the society; and how we can measure the resulting effects". Ethayarajh and Jurafsky [2021] further argue that failures to consider the practical utility of benchmarks has for example made it possible to ignore the discriminatory and environmental damages of AI technologies and allowed for highly energy-inefficient and deeply biased AI models to reach the top of most benchmark leaderboards.
5.4 Narrow Benchmark Diversity and Scope